OpenStudy (anonymous):

Write the radical expression using exponents? 9th root of 3^2

OpenStudy (saifoo.khan):

\[\sqrt[9]{3^{2}}\]

OpenStudy (saifoo.khan):

this?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

\[\sqrt[9]{9}\]

OpenStudy (anonymous):

9th root of 9?

OpenStudy (anonymous):

\[\sqrt[9]{9}\]

OpenStudy (saifoo.khan):

yes!